OpenStudy (anonymous):

one jug holds 7/8 gallons of water. if a faucet can release 16 gallons in 8 minutes. how long would it take to fill one jug with water?

OpenStudy (anonymous):

16 gallons in 8 minutes = 16/8 gallons per minute, or 2 gpm. Lets say the jug could hold 8/8 gallons, or one gallon. It would follow then that it would take 1 gallon/2gallons per minute = 1/2 minutes (the gallon units cancel out) for the jug to fill.

OpenStudy (anonymous):

Then taking the given 7/8 gallon capacity, you would have (7/8gal)/2gpm

OpenStudy (anonymous):

This is a proportion problem. 16 gal : 8 min = 7/8 gal = x min 16/8 = (7/8)/x Cross multiply and solve for x.